OpenStudy (anonymous):

8x=4x-12 whats the value of x A. 3 B.-4 C. 4 D.-3

OpenStudy (anonymous):

subtract 4x from both sides 8x-4x=4x-12-4x 4x=-12

OpenStudy (gabebae):

answer would be -3

OpenStudy (anonymous):

now divide both sides by 4 4x/4=-12/4 x=-3

OpenStudy (anonymous):

Ok thank you.